The times that cars took to refuel at a busy service station are normally distributed with mean 3

minutes and a standard deviation of 0.5 minutes. What is the probability that a car will take

(a) more than 4 minutes to refuel,

(b) not more than 2 minutes to refuel,


1
Expert's answer
2021-11-16T11:39:34-0500

"\mu= 3 \\

\sigma = 0.5"

(a)

"P(X>4) = 1 -P(X<4) \\

= 1 -P(Z< \frac{4-3}{0.5}) \\

= 1 -P(Z< 2) \\

= 1 -0.9772\\

= 0.0228"

(b)

"P(X<2) = P(Z< \frac{2-3}{0.5}) \\

= P(Z< -2) \\

= 0.02275"